OpenStudy (anonymous):

three number which is co-prime to each other are such that the sum product of first two is 551 and that of last two is 1073 what is sum of three number

OpenStudy (raden):

let that numbers are ; a,b,and c 551 = 19*29 and 1073 = 29*73 i think, now u can do it...